OpenStudy (anonymous):

solve each system of equations x+3y=5 3x+4y=20

OpenStudy (anonymous):

solve for what

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

x+3y=5 subtract 3y from each side x=5-3y

OpenStudy (anonymous):

From the first equation, x=5-3y , Plug the value of x from the above into the second equation and solve for y.

OpenStudy (anonymous):

x+3y=5 subtract x from both sides 3y=5-x divide each side by 3 y=\[\frac{ 5-x }{3 }\]

OpenStudy (anonymous):

3x+4y=20 subtract each side by 4y 3x=20-4y divide each side by 3 x=\[\frac{ 20-4y }{ 3 }\]

OpenStudy (anonymous):

3x+4y=20 subtract each side by 3x 4y=20-3x divide each side by 4 y=5-\[\frac{ 3 }{ 4x }\]

OpenStudy (anonymous):

\[3(5-3y) + 4 y = 20 \]\[15-9 y + 4 y = 20 \]\[-5 y = 5 \]y = -1